# Signed Higher-Order Interactions for Brain Disorder Diagnosis via Multi-Channel Transformers

### v3

@@ Abstract
Accurately characterizing the higher-order interactions of brain regions and effectively extracting the interpretable higher-order organizational patterns from Functional Magnetic Resonance Imaging (fMRI) data are crucial for the diagnosis of brain diseases. However, current graph models based on deep learning mainly focus on pairwise patterns, as well as triadic patterns within brain while overlooking more higher-order patterns with signs, limiting an integrated understanding of brain-wide communication. To address these challenges, we propose HOI-Brain (Higher-Order Interaction in Brain Network), a novel computational framework that enables the utilization of signed higher-order interactions and signed organizational patterns in fMRI data for the diagnosis of brain diseases. Specifically, we present a new calculation of co-fluctuations based on Multiplication of Temporal Derivatives to detect higher-order interactions with adequate temporal resolution. Next, we further distinguish positively and negatively synergistic higher-order interactions and encode them in signed weighted simplicial complexes, which can offer detailed insights into the communication within the brain. Moreover, founded upon Persistent Homology theory, two distinct filtration processes are employed in weighted simplicial complexes of the brain to extract signed higher-dimensional neural organisations from a spatiotemporal perspective. Finally, a multi-channel brain Transformer architecture is proposed to holistically integrate information from heterogeneous topological features. Comprehensive experiments across Alzheimer’s disease, Parkinson’s syndrome, and autism spectrum disorder datasets demonstrate the superiority, effectiveness, and interpretability of our framework. The extracted key brain regions and higher-order organizational patterns have been demonstrated to align reliably with domain knowledge from the neuroscience scientific literature, thereby providing direct and meaningful insights. Our code is avaiable.
@@
A novel metric - Multiplication of Temporal Derivatives (MTD) - for quantifying dynamic functional co-fluctuations of group ROIs is proposed.
The first utilization of quadruplet-level interaction signatures and two-dimensional void descriptors extracted by Persistent Homology in the brain enhances diagnostic efficacy.
To the best of our knowledge, this is the first attempt to distinguish between positively and negatively synergistic higher-order interactions.
A novel multi-channel brain network Transformer is proposed to synergistically integrate lower-order edge features with the higher-order topological invariants.
@@ 1 Introduction
In order to extract higher-order topological features that are more readily interpretable for the purpose of brain-disorder diagnosis, recent studies have adopted persistent homology (PH), a topological approach capable of reconstructing HOI structures and delivering state-of-the-art performance in characterising brain topological profiles (Hyekyoung et al., 2011; Sizemore et al., 2018). However, the majority of investigations into higher-order structures in brain networks have focused on 0-dimensional (connected components) and 1-dimensional (cycles) topological profiles formed by nodes and edges. These dimensions do not link HOIs in the brain to higher-order organisations that would provide a truly higher-dimensional perspective (Talesh Jafadideh and Mohammadzadeh Asl, 2022; Bian et al., 2024; Bhattacharya et al., 2025). In order to address the aforementioned limitation, researchers have recently employed synchronisation phenomena to construct more accurate simplicial complexes for modelling HOIs. Furthermore, they have utilised PH to capture 1-dimensional cycles that reflect triplet interactions (Santoro et al., 2024). Comprehensive analyses demonstrate that methods incorporating inferred HOIss among three ROIs outperform traditional pairwise approaches, thereby offering new insights into the higher-order organisation of fMRI time-series data. Nonetheless, the extant evidence indicates that triplet interactions can be decomposed into linear combinations of pairwise interactions, provided that said interactions are linearly decomposable. This finding suggests that some triplets may not in fact represent genuine higher-order phenomena (Delabays et al., 2025; Neuhäuser et al., 2020). Consequently, consideration of HOIs and organisational structures may offer a more profound comprehension of brain function. Furthermore, extant measures of instantaneous co-fluctuation are contingent on extended Pearson correlation, a method that is deficient in temporal resolution when it comes to detecting spatiotemporal interactions among groups of regions (Shine et al., 2015). Finally, the role of signed HOIs, which have the potential to provide valuable diagnostic information in cases of neurological conditions, has been largely overlooked in these studies.
In order to address the methodological challenges identified, a novel computational framework has been proposed. This framework, designated HOI-Brain (Higher-Order Interactions in Brain), facilitates a comprehensive analysis of HOIs in fMRI data 1. Firstly, a novel metric – Multiplication of Temporal Derivatives (MTD) – is introduced with the aim of quantifying dynamic functional co-fluctuations among groups of ROIs. MTD performs element-wise product calculations on the temporal derivatives of blood-oxygen-level-dependent (BOLD) signals, thereby yielding instantaneous co-fluctuation magnitudes for k-node interactions. This approach has been demonstrated to offer a more reliable means of identifying genuine higher-order neural interactions in comparison to Pearson’s correlation. Secondly, adopting a spatiotemporal analytic perspective, we propose a model for brain networks as weighted simplicial complexes. These complexes link instantaneous co-fluctuation of k-node interactions in the brain to the weight of the k-simplex. This facilitates the accommodation of more information about interactions among multiple ROIs in comparison to pairwise networks. The following step involves the extraction of both positive and negative synergistic interactions from the weighted simplicial complexes in order to construct two signed weighted simplicial complexes, respectively. Positively synergistic interactions are indicative of multiple brain regions that exhibit simultaneous activation at a given moment relative to the preceding one, while negatively synergistic interactions indicate that these regions collectively exhibit inhibition at the current moment compared to the prior moment. Collectively, these interactions offer detailed information about complex coordination and communication within the brain, thus improving the effectiveness of brain disease diagnosis. The present study explores the influence of signed HOIs using the self-attention mechanism. The extraction of four types of higher-order topological features is achieved through the implementation of two distinct filtration processes, which are based on Persistent Homology theory. The first type of interaction signature is positive and negative quadruplet-level interaction signatures. These capture irreducible higher-order neural coordination patterns. The second type of interaction signature is positive and negative two-dimensional void descriptors. These characterise the intrinsic geometric organisation of neural activity from a higher-dimensional manifold perspective. Thirdly, given the significance of both lower-order interactions (LOIs) and HOIs in brain network analysis tasks, a multi-channel brain network Transformer is developed to synergistically integrate lower-order edge features with the above four types of higher-order topological invariants. Specifically, higher-order information is injected into a multi-channel Transformer, and an orthonormal clustering readout operation is employed. This operation is based on self-supervised soft clustering and orthonormal projection, and it is used to learn distinguishable cluster-aware lower- and higher-order node embeddings. Subsequently, an attention mechanism is employed to facilitate the adaptive fusion of features encoded by disparate channels. In conclusion, the Multi-Layer Perceptron (MLP) is employed as the prediction head to categorise the various levels of topological features that have been extracted from the multi-channel brain network Transformer.
@@
A novel metric - Multiplication of Temporal Derivatives (MTD) - for quantifying dynamic functional co-fluctuations of group ROIs is proposed. This approach employs element-wise products of temporal derivatives of blood oxygen level-dependent (BOLD) signals to represent instantaneous co-fluctuation magnitudes of k node interactions, which can achieve reliable identification of genuine higher-order neural interactions compared to Pearson correlation.
@@
The first utilization of quadruplet-level interaction signatures and two-dimensional void descriptors in the brain enhances diagnostic efficacy, providing a higher-dimensional perspective for studying higher-order brain structures compared to one-dimensional cycles.
@@
A novel multi-channel brain network Transformer is proposed to synergistically integrate lower-order edge features with the higher-order topological invariants.
@@ 3 Method
The task of diagnosing brain disorders primarily involves inferring specific properties (represented as class labels) from fMRI data in the form of graphs. Given a labeled dataset , where each graph corresponds to a label , the goal is to extract features to train a mapping function that generalizes to unobserved graphs. For example, in studies of Alzheimer’s disease using brain networks, the label space comprises three diagnostic categories: NC (Normal Control), MCI (Mild Cognitive Impairment) and AD (Alzheimer’s Disease). The objective is to extract both lower-order and higher-order topological features from annotated brain network data in order to learn a robust classifier , ensuring its effectiveness on novel, unseen brain connectomes.
3.1 Signed Higher-Order Interaction Representation Strategy
Most traditional network neuroscience approaches construct functional connectivity through pairwise Pearson correlations of fMRI time series (Fornito et al., 2016; Lindquist, 2008). Recent edge-centric frameworks account for interactions of edges by computing instantaneous co-fluctuations through element-wise products of z-scored signals (Faskowitz et al., 2020). Reference (Santoro et al., 2023, 2024) extended this approach to characterize -order interactions when nodes simultaneously deviate from their temporal mean activation at time . Although these techniques based on Pearson correlation have provided insights into the brain’s dynamic HOIs, the methods have generally been limited by the lack of adequate temporal resolution (Shine et al., 2015). Now we give a new way to calculate co-fluctuations between several ROIs.
3.1.1 Calculate Co-fluctuation Through Multiplication of Temporal Derivatives
Considering the above observation, we propose a novel method for estimating instantaneous co-fluctuation magnitudes by incorporating the Multiplication of Temporal Derivatives (Shine et al., 2015). Let us consider an -dimensional real-valued time series with time points, where represents the generic time series recorded from brain region . We first calculate the temporal derivative of each time series by performing a first-order differencing:
for . We normalize each data point by dividing the temporal derivative by its standard deviation , computed over the entire time course. This normalization yields the new time series , expressed as:
where denotes the standard deviation of . Subsequently, the generic element at time of the -scored -order co-fluctuation among new time series is calculated as:
where and represent the time-averaged mean and standard deviation functions, respectively. To differentiate concordant group interactions from discordant ones in a -order product, concordant signs are always positively mapped, while discordant signs are negatively mapped. Formally,
@@ 3.1.1 Calculate Co-fluctuation Through Multiplication of Temporal Derivatives
3.1.2 Construct weighted simplicial complexes
For each temporal instance , we also integrate multi-order co-fluctuations into a weighted simplicial complex . Formally, a -dimensional simplex is defined as a collection of vertices, expressed as . A simplicial complex satisfies the closure condition: every face of a simplex in must also belong to . This structure is enhanced through weight assignments, where numerical values quantify interaction intensities across different simplex dimensions. The resulting framework provides temporal resolution to the evolving patterns of coordinated (and uncoordinated) group-level neural dynamics observed in fMRI signals.
3.1.3 Construct two signed weighted simplicial complexes
It is important to acknowledge that in a weighted simplicial complex, denoted by , there exist two categories of signs for a simplex: concordant signs of positive and discordant signs of negative. The present study focuses exclusively on concordant signs of positive, eschewing the examination of discordant signs of negative. This is due to the potential for such signs to represent a confusing redundancy of information across multiple brain regions (Santoro et al., 2023). The concordant interactions were subsequently categorised into two distinct groups: those that were positively synergistic and those that were negatively synergistic. Positively synergistic interactions are indicative of multiple brain regions that exhibit simultaneous activation at a given moment relative to the preceding one, while negatively synergistic interactions indicate that these regions collectively exhibit inhibition at the current moment compared to the prior moment. Collectively, these interactions offer valuable insights into the complex coordination and communication within the brain, thus improving the effectiveness of brain disease diagnosis. As demonstrated in Fig. 2, it exhibits six distinct signed quadruplet HOIs. For each temporal instance , the weighted simplicial complexes are further extracted to construct two signed weighted simplicial complexes and , namely the positive and negative complexes, respectively.
3.2 Topological Higher-order and Lower-order organization of fMRI signals

### v4

# HOI-Brain: a novel multi-channel transformers framework for brain disorder diagnosis by accurately extracting signed higher-order interactions from fMRI

@@ Abstract
Accurately characterizing the higher-order interactions of brain regions and effectively extracting the interpretable higher-order organizational patterns from functional Magnetic Resonance Imaging (fMRI) data are crucial for the diagnosis of brain diseases. However, current graph models mainly focus on pairwise patterns, as well as triadic patterns within brain while overlooking more higher-order patterns with signs, limiting an integrated understanding of brain-wide communication. To address these challenges, we propose HOI-Brain (Higher-Order Interaction in Brain Network), a novel computational framework that enables the utilization of signed higher-order interactions and signed organizational patterns in fMRI data for the diagnosis of brain diseases. Specifically, we present a new calculation of co-fluctuations based on Multiplication of Temporal Derivatives to detect higher-order interactions with adequate temporal resolution. Next, we further distinguish positively and negatively synergistic higher-order interactions and encode them in the monitonic weighted simplicial complexes, which can offer detailed insights into the communication within the brain. Moreover, we employ Persistent Homology in the monotonic weighted simplicial complexes of the brain to extract signed higher-dimensional neural organisations from a spatiotemporal perspective. Finally, a multi-channel transformers architecture is proposed to holistically integrate information from heterogeneous topological features. Comprehensive experiments across Alzheimer’s disease (AD), Parkinson’s disease (PD), and Autism Spectrum Disorder (ASD) datasets demonstrate the superiority, effectiveness, and interpretability of our framework. The identified key brain regions and higher-order patterns consistently align with neuroscientific evidence, showing concordant HOIs weakening in the brain of AD/ASD but enhancement in PD, offering mechanistic insights into these brain disorders. Our code is avaiable.
@@
We propose the Multiplication of Temporal Derivatives (MTD), a novel metric designed to quantify dynamic functional co-fluctuations across group-level brain regions of interest.
For the first time, we define the monotonic weighted simplicial complex for brain networks, extracting quadruplet-level interaction signatures and two-dimensional void descriptors via Persistent Homology.
We propose the first to discriminate positive and negative synergistic higher-order interactions and reveal their disease-specific associations in brain disorders.
We introduce a novel multi-channel brain network Transformer to synergistically integrate lower-order edge features with the higher-order topological invariants.
We identify disease-stage-dependent higher-order organizational patterns, showing gradually weakened concordant higher-order interactions (HOIs) from healthy controls to AD/ASD patients, while revealing an opposite trend in Parkinson’s disease.
@@ 1 Introduction
In order to extract higher-order topological features that are more readily interpretable for the purpose of brain-disorder diagnosis, recent studies have adopted Persistent Homology, a topological approach capable of reconstructing HOI structures and delivering state-of-the-art performance in characterising brain topological profiles (Hyekyoung et al., 2011; Sizemore et al., 2018). However, the majority of investigations into higher-order structures in brain networks have focused on 0-dimensional (connected components) and 1-dimensional (cycles) topological profiles formed by nodes and edges. These dimensions do not link HOIs in the brain to higher-order organisations that would provide a truly higher-dimensional perspective (Talesh Jafadideh and Mohammadzadeh Asl, 2022; Bian et al., 2024; Bhattacharya et al., 2025). In order to address the aforementioned limitation, researchers have recently employed synchronisation phenomena to construct more accurate simplicial complexes for modelling HOIs. Furthermore, they have utilised Persistent Homology to capture 1-dimensional cycles that reflect triplet interactions (Santoro et al., 2024). Comprehensive analyses demonstrate that methods incorporating inferred HOIss among three ROIs outperform traditional pairwise approaches, thereby offering new insights into the higher-order organisation of fMRI time-series data. Nonetheless, the extant evidence indicates that triplet interactions can be decomposed into linear combinations of pairwise interactions, provided that said interactions are linearly decomposable. This finding suggests that some triplets may not in fact represent genuine higher-order phenomena (Delabays et al., 2025; Neuhäuser et al., 2020). Consequently, consideration of HOIs and organisational structures may offer a more profound comprehension of brain function. Furthermore, extant measures of instantaneous co-fluctuation are contingent on extended Pearson correlation, a method that is deficient in temporal resolution when it comes to detecting spatiotemporal interactions among groups of regions (Shine et al., 2015). Finally, the role of signed HOIs, which have the potential to provide valuable diagnostic information in cases of neurological conditions, has been largely overlooked in these studies.
In order to address the methodological challenges identified, a novel computational framework has been proposed in this paper. As shown in Figure 1, this framework, designated HOI-Brain (Higher-Order Interactions in Brain), facilitates a comprehensive analysis of HOIs in fMRI data.
Firstly, a novel metric – Multiplication of Temporal Derivatives (MTD) – is introduced with the aim of quantifying dynamic functional co-fluctuations among groups of ROIs. MTD performs element-wise product calculations on the temporal derivatives of blood-oxygen-level-dependent (BOLD) signals, thereby yielding instantaneous co-fluctuation magnitudes for k-node interactions. This approach will offer a more reliable means of identifying genuine higher-order neural interactions in comparison to Pearson’s correlation.
Secondly, we further divided concordant interactions into two distinct groups: those exhibiting positive synergy and those exhibiting negative synergy. Positively synergistic interactions are indicative of multiple brain regions that exhibit simultaneous activation at a given moment relative to the preceding one, while negatively synergistic interactions indicate that these regions collectively exhibit inhibition at the current moment compared to the prior moment. Collectively, these interactions offer detailed information about complex coordination and communication within the brain, thus improving the effectiveness of brain disease diagnosis. Our study explores the influence of signed HOIs using the self-attention mechanism.
Thirdly, adopting a spatiotemporal analytic perspective, we encode these signed co-fluctuations into the monotonic weighted simplicial complexes. These complexes link instantaneous co-fluctuation of k-node interactions in the brain to the weight of the k-simplex. This facilitates the accommodation of more information about interactions among multiple ROIs in comparison to pairwise networks. The extraction of four types signed of higher-order topological which from these monotonic weighted simplicial complexes are based on Persistent Homology theory. The first type of interaction signature is positive and negative quadruplet-level interaction signatures. These capture irreducible higher-order neural coordination patterns. The second type of interaction signature is positive and negative two-dimensional void descriptors. These characterise the intrinsic geometric organisation of neural activity from a higher-dimensional manifold perspective.
Then, given the significance of both lower-order interactions (LOIs) and HOIs in brain network analysis tasks, a multi-channel brain network Transformer is developed to synergistically integrate lower-order edge features with the above four types of higher-order topological invariants. Specifically, higher-order information is injected into a multi-channel Transformer, and an orthonormal clustering readout operation is employed. This operation is based on self-supervised soft clustering and orthonormal projection, and it is used to learn distinguishable cluster-aware lower- and higher-order node embeddings. Subsequently, an attention mechanism is employed to facilitate the adaptive fusion of features encoded by disparate channels. Finally, the Multi-Layer Perceptron (MLP) is employed as the prediction head to categorise the various levels of topological features that have been extracted from the multi-channel brain network Transformer.
@@
We propose a novel metric - Multiplication of Temporal Derivatives (MTD) - for quantifying dynamic functional co-fluctuations of group ROIs. This approach employs element-wise products of temporal derivatives of blood oxygen level-dependent (BOLD) signals to represent instantaneous co-fluctuation magnitudes of k-node interactions, which can achieve reliable identification of genuine higher-order neural interactions compared to Pearson correlation.
@@
For the first time, we define the monotonic weighted simplicial complex for brain networks, extracting quadruplet-level interaction signatures and two-dimensional void descriptors via persistent homology. This provide a higher-dimensional perspective for studying higher-order brain structures compared to one-dimensional cycles.
@@
We propose a novel multi-channel brain network Transformer to synergistically integrate lower-order edge features with the higher-order topological invariants.
5.
We identify disease-stage-dependent higher-order organizational patterns, showing gradually weakened concordant HOIs from healthy controls to AD/ASD patients, while revealing an opposite trend in Parkinson’s disease.
@@ 2.2 Persistent Homology on Brain Connectome
3 Preliminaries
In this section, we firstly present the problem formulation of this paper. Then, we propose a general method for mining HOIs in time series data using Multiplication of Temporal Derivatives. Finally, we introduce a new definition of monotonic weighted simplicial complex. These will provide the theoretical basis for the methods of brain disease diagnosis in the following text.
3.1 Problem Formulation
The task of diagnosing brain disorders primarily involves inferring specific properties (represented as class labels) from fMRI data in the form of graphs. Given a labeled dataset , where each graph corresponds to a label , the goal is to extract features to train a mapping function that generalizes to unobserved graphs. For example, in studies of Alzheimer’s disease using brain networks, the label space comprises three diagnostic categories: CN (Cognitively Normal), MCI (Mild Cognitive Impairment) and AD (Alzheimer’s Disease). The objective is to extract both lower-order and higher-order topological features from annotated brain network data in order to learn a robust classifier , ensuring its effectiveness on novel, unseen brain connectomes.
3.2 Multiplication of Temporal Derivatives for Processing Multivariate Time Series
To better capture HOIs in multivariate time series data with adequate temporal resolution, we propose a novel method for estimating instantaneous co-fluctuation magnitudes between several variables.
Let represent the generic time series recorded from variable . We first calculate the temporal derivative of each time series by performing a first-order differencing:
for . Then we normalize each data point by dividing the temporal derivative by its standard deviation , computed over the entire time course. This normalization yields the new time series as:
Then the new representing vector for the ith variable is denoted as .
Subsequently, the generic element at time of the -scored -order co-fluctuation among new time series is calculated as:
where and represent the time-averaged mean and standard deviation functions, respectively; denotes the product of the elements; represents the element-wise product of vectors.
To differentiate concordant group interactions from discordant ones in a -order product, concordant signs are always positively mapped, while discordant signs are negatively mapped. Formally,
It is noteworthy that our proposed method, by computing the product of first-order difference sequences of multivariate time series, enables the investigation of higher-order dynamic interactions, thereby providing a viable approach for exploring complex nonlinear coupling mechanisms in multivariate time series data.
3.3 The Monotonic Weighted Simplicial Complex
Definition 1.
(Simplicial Complex) An abstract simplicial complex on a finite vertex set is a collection of some non-empty subsets of satisfying:
1.
For every , the singleton belongs to .
2.
If and is a non-empty subset, then .
Each element is called a simplex. A simplex with vertices is a -simplex, representing a -order interaction. A subset satisfying is called a face of .
Previous studies assigned arbitrary weights to simplices without imposing closure conditions on the weights (Baccini et al., 2022). To explore some special higher-order structures in the simplicial complex, here we propose a new definition of weighted simplicial complexes.
Definition 2.
(Monotonic Weighted Simplicial Complex) Let be an abstract simplicial complex on vertex set , where each simplex is a non-empty finite subset of . A monotonic weighted simplicial complex is a pair with a weight function satisfying the following lower closure condition:
1.
(Lower closure) For any , any face , .
In addition, a monotonic weighted simplicial complex is also a pair with a weight function satisfying the following upper closure condition:
1.
(Upper closure) For any , any face , .
As shown in Figure 2, it shows two types of monotonic weighted simplicial complexes. The new definition allows us to explore some special higher-order structures satisfying the monotonicity closure conditions in the weighted simplicial complex, providing new insights.
@@ 4 Method
In this section, we introduce the method of HOI-Brain for diagnosing brain diseases, which captures signed high-order interactions and signed higher-order organizational patterns in fMRI data. HOI-Brain consists of four parts: calculating signed high-order interactions through Multiplication of Temporal Derivatives, constructing monotonic weighted simplicial complexes, extracting higher-order and low-order organizations in fMRI data, and classifying with a multi-channel brain network Transformer.
4.1 Calculate Signed Higher-Order Interactions Through Multiplication of Temporal Derivatives.
Recently, some computational methods have focused on capturing dynamic higher-order interactions in the brain (Faskowitz et al., 2020; Santoro et al., 2023, 2024), but these methods based on extended Pearson correlation have generally been limited by the lack of adequate temporal resolution, especially when capturing more higher-order interactions, such as quadruplets (Shine et al., 2015). To better capture higher-order interactions in the brain, we utilize the designed Multiplication of Temporal Derivatives on fMRI data, which is described in Section 3.2.
In particular, individual fMRI data are transformed into original fMRI signals through a preprocessing pipeline. Temporal derivatives of each time series (length ) are computed via first-order differencing, followed by normalization where each derivative is divided by its standard deviation across the entire time course. All possible -order time series are generated by element-wise products of normalized time series. These series are further z-scored to ensure cross--order comparability. These -order time series represent instantaneous co-fluctuation magnitudes of -node interactions (e.g., edges, triangles, quadruplets).
@@
It is important to acknowledge that for all -order co-fluctuations, there exist two categories of signs for co-fluctuations: concordant signs and discordant signs, as shown in Figure 3. This study focuses exclusively on concordant signs, eschewing the examination of discordant signs. This is due to the potential for discordant signs to represent a confusing redundancy of information across multiple brain regions (Santoro et al., 2023). The concordant interactions were subsequently categorised into two distinct groups:
1) Positively synergistic interactions are indicative of multiple brain regions that exhibit simultaneous activation at a given moment relative to the preceding one. These interactions are mathematically defined by the conditions:
where denotes the weight assigned to the -simplex , denotes the first-order temporal difference of the activity value at node . Let denote the set of all k-order positively synergistic interactions (), with their weights .
2) Negatively synergistic interactions indicate that these regions collectively exhibit inhibition at the current moment compared to the prior moment. These interactions are mathematically defined by the conditions:
where denotes the weight assigned to the -simplex , denotes the first-order temporal difference of the activity value at node . Let denote the set of all k-order negatively synergistic interactions (), with their weights .
Collectively, distinguishing these interactions can offer valuable insights into the complex coordination and communication within the brain, thus improving the effectiveness of brain disease diagnosis.
4.2 Construct Monotonic Weighted Simplicial Complexes
Recent advances in neuroimaging analysis have demonstrated the utility of representing brain as simplicial complexes (Sizemore et al., 2018; Santoro et al., 2023), which preserves the inherent higher-order dependencies in neural systems that are critical for understanding cognitive processes and pathological mechanisms compared to two-way networks (Barrat et al., 2004). Based on this, we further propose the concept of monotonic weighted simplicial complex in Section 3.3 to explore some special higher-order structures in the brain and model neurobiological principles where functional coherence depends on hierarchical organization.
Specifically, for each temporal instance , we can associate the k-order co-fluctuations with k-simplexes to construct four different monotonic weighted simplicial complexes. For example, the monotonic weighted simplicial complex that satisfies the lower closure condition and is linked with positively synergistic interactions. The other three monotonic weighted simplicial complexes can be analogized from this. Specifically, in these simplicial complexes, a 0-simplex is a node associated with a ROI, a 1-simplex is a edge associated with pairwise interaction, a 2-simplex is a triangle associated with triplet interaction, a 3-simplex is a quadruplet associated with quadruplet interaction.
To construct these monotonic weighted simplicial complexes, we leverage the concept of filtration from Persistent Homology theory (Talesh Jafadideh and Mohammadzadeh Asl, 2022), a computational topology technique adept at analyzing high-dimensional datasets. Specifically, we design two types of filtrations to generate the monotonic weighted simplicial complexes that satisfies the lower closure condition, and the monotonic weighted simplicial complexes that satisfies the upper closure condition.
The ascending filtration: Sort the weights in the set of all positively synergistic simplices in ascending order and use the parameter to scan this sequence . At each step , we include all 1-simplices, 2-simplices, and 3-simplices that satisfy the lower closure condition of the monotonic weighted simplicial complex to generate a monotonic weighted simplicial complex . The set of all negatively synergistic simplices is processed in the same way to generate a monotonic weighted simplicial complex . These are defined as:
The descending filtration: Sort the weights in the set of all positively synergistic simplices in descending order and use the parameter to scan this sequence . At each step , we include all 1-simplices, 2-simplices, and 3-simplices that satisfy the upper closure condition of the monotonic weighted simplicial complex to generate a monotonic weighted simplicial complexes . The set of all negatively synergistic simplices is processed in the same way to generate a weighted simplicial complexes . These are defined as:
These weighted simplicial complexes encode the brain from different perspectives, providing the potential to further explore higher-order organizational patterns in the brain.
4.3 Topological Lower-Order and Higher-Order Organizations of fMRI Signals